# From Synthetic to Real: Unveiling the Power of Synthetic Data for Video Person Re-ID

### v2

@@ Abstract
In this study, we investigate the novel challenge of cross-domain video-based person re-identification (Re-ID). Here, we utilize synthetic video datasets as the source domain for training and real-world videos for testing, notably reducing the reliance on expensive real data acquisition and annotation. To harness the potential of synthetic data, we first propose a self-supervised domain-invariant feature learning strategy for both static and dynamic (temporal) features. Additionally, to enhance person identification accuracy in the target domain, we propose a mean-teacher scheme incorporating a self-supervised ID consistency loss. Experimental results across five real datasets validate the rationale behind cross-synthetic-real domain adaptation and demonstrate the efficacy of our method. Notably, the discovery that synthetic data outperforms real data in the cross-domain scenario is a surprising outcome. The code and data will be publicly available at https://github.com/XiangqunZhang/UDA_Video_ReID.
Keywords:
video-based person re-identification cross domain synthetic videos
journal: International Journal of Computer Vision
1 Introduction
Person re-identification (Re-ID) stands as a classical problem in computer vision, experiencing rapid evolution in recent years. This task holds significant real-world utility in areas like video surveillance and criminal investigation. We first review the development of person Re-ID as shown in Fig. 1. On one hand, in terms of the data format, early works mainly use a single image for retrieving the specific person (Li et al., 2014; Ahmed et al., 2015). Later, the sequential video data are applied to person Re-ID since it contains more plentiful information (You et al., 2016; Ouyang et al., 2019; Chen et al., 2020a), e.g., the human gait. On the other hand, in terms of the dataset conditions, early works focus on the data distributions in the training and testing datasets are similar. In recent works, unsupervised domain adaptive (UDA) person Re-ID is proposed aiming to train the model on the source domain and test it on the target domain. This is more practical since the characteristics of data in the training and testing sets are always various in the real-world application, including the illumination, viewpoint, etc. For UDA person Re-ID, several works (Zhang et al., 2021; Sun and Zheng, 2019) apply the synthetic data for model training to alleviate the burden of data collection and annotation.
@@ 1 Introduction
  • Extensive experimental results demonstrate the effectiveness of the proposed method, which outperforms the comparative methods over 5.1%, 11.2%, 11.3%, 2.9%, and 4.1% on five real evaluation datasets. Further analysis also verifies that the utilization of synthetic videos for cross-domain video person Re-ID is effective, even better than real videos.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related works. Section 3 elaborates on the proposed approach in detail. Section 4 provides the experimental results and analysis. After that, we provide a brief conclusion including the future work in Section 5.
2 Related Work
@@
Video-based person Re-ID is different from image-based Re-ID, which relies on videos as input and contains richer spatial-temporal information compared to static images. Therefore, exploring spatial-temporal cues is crucial in video-based person Re-ID problems. With respect to the stable spatial representation, the method in Gu et al. (2020) builds the appearance-preserving module to solve the problem of appearance destruction. Another work (Bai et al., 2022) combines salient-to-broad attention and interaction-driven message passing to capture spatial information. Specific to the challenge of changing clothes, CCVID (Gu et al., 2022) proposes a clothes-based adversarial loss to mine clothes-irrelevant features for clothes-changing video person Re-ID. Similarly, to obtain more effective temporal representations, many technologies are developed, such as complementary features of successive frames (Hou et al., 2020), temporal reciprocal learning for temporal information accumulation (Liu et al., 2021b), and temporal shift attention module (Wu et al., 2022a). Furthermore, some approaches focus on aggregating both spatial and temporal features to address spatial and temporal distractors. A spatial-temporal aggregation module is proposed to enhance the aggregation capability (Wang et al., 2021). Another work (Eom et al., 2021) leverages spatial-temporal memories to refine frame-level person representations to get sequence-level representation. The axial-attention strategy is utilized to capture spatial and temporal relations for addressing the misalignment issue in video-based Re-ID (Liu et al., 2021a). These methods aim to fuse features from multiple temporal and spatial scales to obtain comprehensive representations. We find that although significant progress has been made in video person Re-ID, in which cross-domain scenarios have been relatively limited explored, not to mention that across the synthetic and real domains.
@@
Unsupervised domain adaptive (UDA) person Re-ID relies on transferring knowledge learned from the source domain to the target domain. First, some approaches enhance the accuracy of representation learning in the target domain by improving clustering methods. The method in Zheng et al. (2021a) utilizes online clustering and label optimization to improve the accuracy of representation learning in the target domain. Another work (Hu et al., 2022) considers the multi-camera distribution and temporal continuity in re-ID to effectively reduce clustering errors. A density-based clustering with adaptive sample augmentation and discriminative feature learning is utilized to estimate and augment person clusters in target domains (Zhai et al., 2020). Zheng et al. (2021b) dynamically and progressively update the pseudo labels based on hierarchical cluster dynamics. Second, some methods focus on generating and refining pseudo-labels to mitigate the impact of label noise in the target domain. Wu et al. (2022b) propose a multi-centroid memory module that captures different ID information within a cluster. Han et al. (2022) propose a probabilistic uncertainty-guided progressive label refinement approach. The work in Li et al. (2022) utilizes a mean teacher model for online label refining. Cho et al. (2022) leverage both global and local context information to mitigate label noise. Third, some methods reduce the domain offset between the source domain and the target domain by domain distribution alignment, e.g., the graph convolutional network-based multi-domain information fusion module (Bai et al., 2021), and feature spaces constructing and alignment with the soft labels (Yu et al., 2019). The advancements in UDA methods have significantly contributed to cross-domain person Re-ID. However, these methods predominantly focus on image-based tasks and have not yet explored video-based tasks. Differently, this work tackles the challenging task of cross-domain Re-ID in the context of video data, aiming to learn robust and invariant person representations during the temporal variations.
@@
Person Re-ID with synthetic data. In recent years, the difficulty of labeling real data has made synthetic data shine in the field of person Re-ID. The work in Bak et al. (2018) proposes to synthesize realistic images and minimize the domain shift between synthetic and real images. Later, a number of large-scale image-based synthetic datasets (Sun and Zheng, 2019; Zhang et al., 2021; Wang et al., 2022b) have been proposed for person Re-ID, which provide realistic appearance and clothing by generating virtual human models as a cost-effective alternative to collecting and labeling real data. A series of methods (Xiang et al., 2022, 2020; Kang, 2022) for cross-domain person Re-ID have demonstrated the effectiveness of cross-domain design from synthetic to real. A recent work (Wang et al., 2022a) presents a large-scale video-based synthetic person Re-ID dataset that demonstrates the potential of synthetic data for person Re-ID in challenging scenarios. As discussed above, research on synthetic data-based person Re-ID mainly focuses on image-based tasks, the exploration of which using video data is limited and worth studying.
3 Proposed Method
3.1 Problem formulation
@@
3.2 Multi-level domain invariant feature learning
Basic feature extraction. Given the videos $\mathcal{V}^{\mathrm{L}}$ in the source domain, we first apply an existing deep learning model, e.g., SINet (Bai et al., 2022), as the backbone for video Re-ID feature extraction
@@ 3.2 Multi-level domain invariant feature learning
where $\mathbf{f}_{i}$ denotes the feature of input sequence, $\mathbf{d}_{i}$ denotes the binary domain prediction result of video $\mathbf{f}_{i}$, and $\mathrm{Cla}_{v}$ also denotes the binary classification. The domain label of each frame or video is easy to obtain; hence the above training can be conducted in a self-supervised manner. Besides, we further stitch the features from multiple sequences with a random permutation, as shown in Fig. 2. The number of sequences for combination is arbitrary, e.g., 2,4,8, etc. Then, we aim to predict the domain distribution of the combined features. The prediction can not be achieved by a classification problem since the number of sequences in each combined feature is various. This way, we design a regression strategy for the domain prediction. Specifically, for $V$ sequence features $\mathbf{f}_{i}$, we use the temporal feature aggregation method (Chen et al., 2020b) to form the combined feature as $\mathbf{F}_{g}$, where $g$ denotes the index. Then we predict its domain distribution by
@@
Note that, during training, we apply the adversarial training strategy by using a gradient reversal layer (GRL) (Ganin and Lempitsky, 2015), which reverses the gradient signs during back-propagation. The basic insight is that, with GRL, the domain discrimination ability gets stronger as the consistency of cross-domain features increases. This way, the classification will be optimized to gradually align the feature distributions between the two domains. We find that the method is effective in this task since the synthetic and real domains are more discriminative than those with both real source and target domains, which boosts the domain invariant feature learning. By integrating these three auxiliary tasks, we achieve the multi-granularity domain-invariant feature learning from the single frame, single video, and stitched multiple videos. This enables us to handle the invariant feature learning under cross-domain and over-time variations effectively.
3.3 Consistency learning on unlabeled real data
@@
where ${\bf{f}}^{\mathcal{S}}_{j}$ and ${\bf{f}}^{\mathcal{T}}_{j}$ denote the feature through the student and teacher networks, respectively. A straightforward idea to estimate the consistency between $\mathcal{T}$ and $\mathcal{S}$ is to calculate the distance between ${\bf{f}}^{\mathcal{S}}_{j}$ and ${\bf{f}}^{\mathcal{T}}_{j}$. This may lead to the gradient computation issue because of the large feature distance. As discussed in previous works (Chen et al., 2020b), a better manner to estimate the consistency is to use the result of the target task rather than the middle features. This way, a plausible approach is to apply the classification based on the features to obtain the person IDs. However, this brings two problems. First, the number of IDs in dataset $\mathcal{V}^{\rm U}$ is not provided in prior, i.e., the number of classification categories is unknown. Second, the ID indexes obtained from the classification of $\mathcal{T}$ and $\mathcal{S}$ are unaligned. To address the above problems, we propose a clustering-based identification consistency estimation strategy.
@@ 3.4 Implementation details
which are controlled by a learnable weight vector to determine the ratio among them. The student network is first trained with the supervised losses $\mathcal{L}_{\rm cse},\mathcal{L}_{\rm tri}$ and the domain discriminative loss $\mathcal{L}^{\rm S}$. For the training using the mean teacher strategy, the weights of the teacher network are initialized with those of the student network. The EMA method is used to update the weights of the teacher network in each training iteration. Then, the consistency loss $\mathcal{L}^{\mathrm{C}}_{1},\mathcal{L}^{\mathrm{C}}_{2}$ can be applied.
@@
We use the K-means algorithm to partition the video-level features into $M$ clusters, where $M$ is set as 16. We use 4 $\times$ NVIDIA RTX 2080ti for model training and apply the Adam (Kingma and Ba, 2014) algorithm for optimization. During the testing phase, given a query, the trained student network retrieves a ranked list from the gallery.
4 Experimental Results
4.1 Setup
@@
Synthetic video Re-ID dataset. Previous work (Wang et al., 2022a) uses the highly realistic video game Grand Theft Auto (GTA) V to construct a large-scale dataset for clothes-changing video person Re-ID. The videos were collected in a surveillance scenario, where 10 monitoring cameras were installed across five different scenes. We notice that this dataset mainly focuses on the clothes-changing challenge in the Re-ID problem, in which the persons have various (2-37) suits. Based on this dataset, we built a Synthetic-domain-adaptive Video person Re-ID dataset, namely SVReID. Specifically, we first only consider the non-clothing change scenario. From the training set in  Wang et al. (2022a), we extract a subset by selecting the highest number of sequences for each person without clothing variation and get SVReID including 1,164 sequences from 333 identities. We also consider the clothes-changing problem in some real-world video Re-ID datasets, e.g., CCVID (Gu et al., 2022). This way, we further extend SVReID by including the clothes-changing videos, i.e., we select the top-8 highest number of sequences with clothing variations and get a total of 3,187 sequences, which form SVReID (w cc). Note that this work mainly focuses on the cross-domain problem, thus we keep the clothing scenario of the training data the same as the corresponding testing dataset. This prevents the model from excessively focusing on irrelevant data, and we can compare the performance of different methods under the same scenario. This synthetic dataset has two main advantages. Reality: The videos generated by GTA-V are very realistic, as shown in Fig. 3, the involved persons have various gaits, silhouettes even changed clothes. Diversity: It offers a wider range of various conditions, such as the camera view angles, illumination, and occlusion/clutter scenarios, which can not be easily achieved in real data.
Real-world video Re-ID datasets. In this work, we evaluate the effectiveness of our domain-adaptive Re-ID method on four prominent real-world video datasets, i.e., MARS (Zheng et al., 2016), iLIDS-VID (Wang et al., 2014), PRID-2011 (Hirzer et al., 2011), LS-VID (Li et al., 2019), and CCVID (Gu et al., 2022).
@@
$\bullet$ CCVID is a Re-ID dataset CCVID derived from a gait dataset FVG (Zhang et al., 2019). The characteristic of it is with clothing change setting. CCVID contains 2,856 video sequences from 226 identities, with an average of 2 outfits per person, and an average of 121 frames per sequence.
@@
Table 1: Comparison with the state-of-the-art Re-ID methods on four pairs of cross-domain datasets, i.e., from the synthetic SVReID to the real datasets including Mars, iLIDS-VID, PRID, and LS-VID. (%)

### v3

# Synthetic-To-Real Video Person Re-ID

@@ Abstract
Person re-identification (Re-ID) is an important task and has significant applications for public security and information forensics, which has progressed rapidly with the development of deep learning. In this work, we investigate a novel and challenging setting of Re-ID, i.e., cross-domain video-based person Re-ID. Specifically, we utilize synthetic video datasets as the source domain for training and real-world videos for testing, notably reducing the reliance on expensive real data acquisition and annotation. To harness the potential of synthetic data, we first propose a self-supervised domain-invariant feature learning strategy for both static and dynamic (temporal) features. Additionally, to enhance person identification accuracy in the target domain, we propose a mean-teacher scheme incorporating a self-supervised ID consistency loss. Experimental results across five real datasets validate the rationale behind cross-synthetic-real domain adaptation and demonstrate the efficacy of our method. Notably, the discovery that synthetic data outperforms real data in the cross-domain scenario is a surprising outcome. The code and data are publicly available at https://github.com/XiangqunZhang/UDA_Video_ReID.
Index Terms:
Video-based person re-identification, Cross domain, Synthetic videos.
I Introduction
Person re-identification (Re-ID) stands as a classical problem in artificial intelligence, which holds significant real-world applications for public security and information forensics, e.g., intelligent video surveillance and criminal investigation [1, 2]. This task experiences rapid evolution in recent years. We first review the development of person Re-ID as shown in Fig. 1. On one hand, in terms of the data format, early works mainly use a single image for retrieving the specific person [3, 4]. Later, the sequential video data are applied to person Re-ID since it contains more plentiful information [5, 6, 7], e.g., the human gait. On the other hand, in terms of the dataset conditions, early works focus on the data distributions in the training and testing datasets are similar. In recent works, unsupervised domain adaptive (UDA) person Re-ID is proposed aiming to train the model on the labeled source domain and unlabeled target domain data, and test it on the target domain. This is more practical since the characteristics of data in the training and testing sets are always various in the real-world application, including the illumination, viewpoint, etc. For UDA person Re-ID, several works [8, 9] apply the synthetic data for model training to alleviate the burden of data collection and annotation.
@@ I Introduction
  • Extensive experimental results demonstrate the effectiveness of the proposed method, which outperforms the comparative methods over 5.1%, 11.2%, 11.3%, 2.9%, and 4.1% mAP score on five real evaluation datasets. Further analysis also verifies that the utilization of synthetic videos for cross-domain video person Re-ID is effective, even better than real videos.
@@
The remainder of this paper is organized as follows. Section II reviews the related works. Section III elaborates on the proposed approach in detail. Section IV provides the experimental results and analysis. After that, we provide a brief conclusion including the future work in Section V.
II Related Work
@@
Video-based person Re-ID is different from image-based Re-ID, which relies on videos as input and contains richer spatial-temporal information compared to static images. Therefore, exploring spatial-temporal cues is crucial in video-based person Re-ID problems. With respect to the stable spatial representation, the method in [10] builds the appearance-preserving module to solve the problem of appearance destruction. Another work [11] combines salient-to-broad attention and interaction-driven message passing to capture spatial information. Specific to the challenge of changing clothes, CCVID [12] proposes a clothes-based adversarial loss to mine clothes-irrelevant features for clothes-changing video person Re-ID. Similarly, to obtain more effective temporal representations, many technologies are developed, such as complementary features of successive frames [13], temporal reciprocal learning for temporal information accumulation [14], and temporal shift attention module [15]. Furthermore, some approaches focus on aggregating both spatial and temporal features to address spatial and temporal distractors. A spatial-temporal aggregation module is proposed to enhance the aggregation capability [16]. Another work [17] leverages spatial-temporal memories to refine frame-level person representations to get sequence-level representation. The axial-attention strategy is utilized to capture spatial and temporal relations for addressing the misalignment issue in video-based Re-ID [18]. DCCT [19] develops a deeply-coupled convolution-transformer with spatial-temporal complementary learning for video-based person Re-ID. These methods aim to fuse features from multiple temporal and spatial scales to obtain comprehensive representations. More recently, the multi-modal pre-trained models are also used to the Re-ID task, such as TF-CLIP [20]. We find that although significant progress has been made in video person Re-ID, in which cross-domain scenarios have been relatively limited explored, not to mention that across the synthetic and real domains.
@@
Unsupervised domain adaptive (UDA) image-based person Re-ID relies on transferring knowledge learned from the source domain to the target domain. First, some approaches enhance the accuracy of representation learning in the target domain by improving clustering methods. The method in [21] utilizes online clustering and label optimization to improve the accuracy of representation learning in the target domain. Another work [22] considers the multi-camera distribution and temporal continuity in re-ID to effectively reduce clustering errors. A density-based clustering with adaptive sample augmentation and discriminative feature learning is utilized to estimate and augment person clusters in target domains [23]. Another work [24] dynamically and progressively updates the pseudo labels based on hierarchical cluster dynamics. Liao et al. propose an efficient mini-batch sampling method, called graph sampling (GS), for large-scale deep metric learning [25]. Second, some methods focus on generating and refining pseudo-labels to mitigate the impact of label noise in the target domain. The authors of [26] propose a multi-centroid memory module that captures different ID information within a cluster. Han et al. [27] propose a probabilistic uncertainty-guided progressive label refinement approach. The work in [28] utilizes a mean teacher model for online label refining. Another work [29] leverages both global and local context information to mitigate label noise. Third, some methods reduce the domain offset between the source domain and the target domain by domain distribution alignment, e.g., the graph convolutional network-based multi-domain information fusion module [30], and feature spaces constructing and alignment with the soft labels [31], the extraction of a support set composed of source images that maximizes the similarity with the target data [32]. The advancements in UDA methods have significantly contributed to cross-domain person Re-ID. However, these methods predominantly focus on image-based tasks and have not yet explored video-based tasks. Differently, this work tackles the challenging task of cross-domain Re-ID in the context of video data, aiming to learn robust and invariant person representations during the temporal variations.
@@
Person Re-ID with synthetic data. In recent years, the difficulty of labeling real data has made synthetic data shine in the field of person Re-ID [33]. The work in [34] proposes to synthesize realistic images and minimize the domain shift between synthetic and real images. Later, a number of large-scale image-based synthetic datasets [9, 35, 8, 36] have been proposed for person Re-ID, which provide realistic appearance and clothing by generating virtual human models as a cost-effective alternative to collecting and labeling real data. More recently, the diffusion based method is also applied to the synthetic Re-ID data generation  [37]. A series of methods [38, 39, 40] for cross-domain person Re-ID have demonstrated the effectiveness of cross-domain design from synthetic to real. However, as discussed above, research on synthetic data-based person Re-ID mainly focuses on image-based tasks, the exploration of which using video data is limited and worth studying. A recent work [41] presents a large-scale video-based synthetic person Re-ID dataset that demonstrates the potential of synthetic data for person Re-ID in challenging scenarios, which is available for the synthetic-to-real video Re-ID task in this work.
III Proposed Method
III-A Problem formulation
@@
III-B Multi-level domain invariant feature learning
Basic feature extraction. Given the videos $\mathcal{V}^{\mathrm{L}}$ in the source domain, we first apply an existing deep learning model, e.g., SINet [11], as the backbone for video Re-ID feature extraction
@@ III-B Multi-level domain invariant feature learning
where $\mathbf{f}_{i}$ denotes the feature of input sequence, $\mathbf{d}_{i}$ denotes the binary domain prediction result of video $\mathbf{f}_{i}$, and $\mathrm{Cla}_{v}$ also denotes the binary classification. The domain label of each frame or video is easy to obtain; hence the above training can be conducted in a self-supervised manner. Besides, we further stitch the features from multiple sequences with a random permutation, as shown in Fig. 2. The number of sequences for combination is arbitrary, e.g., 2,4,8, etc. Then, we aim to predict the domain distribution of the combined features. The prediction can not be achieved by a classification problem since the number of sequences in each combined feature is various. This way, we design a regression strategy for the domain prediction. Specifically, for $V$ sequence features $\mathbf{f}_{i}$, we use the temporal feature aggregation method [42] to form the combined feature as $\mathbf{F}_{g}$, where $g$ denotes the index. Then we predict its domain distribution by
@@
Note that, during training, we apply the adversarial training strategy by using a gradient reversal layer (GRL) [43], which reverses the gradient signs during back-propagation. The basic insight is that, with GRL, the domain discrimination ability gets stronger as the consistency of cross-domain features increases. This way, the classification will be optimized to gradually align the feature distributions between the two domains. We find that the method is effective in this task since the synthetic and real domains are more discriminative than those with both real source and target domains, which boosts the domain invariant feature learning. By integrating these three auxiliary tasks, we achieve the multi-granularity domain-invariant feature learning from the single frame, single video, and stitched multiple videos. This enables us to handle the invariant feature learning under cross-domain and over-time variations effectively.
III-C Consistency learning on unlabeled real data
@@
where ${\bf{f}}^{\mathcal{S}}_{j}$ and ${\bf{f}}^{\mathcal{T}}_{j}$ denote the feature through the student and teacher networks, respectively. A straightforward idea to estimate the consistency between $\mathcal{T}$ and $\mathcal{S}$ is to calculate the distance between ${\bf{f}}^{\mathcal{S}}_{j}$ and ${\bf{f}}^{\mathcal{T}}_{j}$. This may lead to the gradient computation issue because of the large feature distance. As discussed in previous works [42], a better manner to estimate the consistency is to use the result of the target task rather than the middle features. This way, a plausible approach is to apply the classification based on the features to obtain the person IDs. However, this brings two problems. First, the number of IDs in dataset $\mathcal{V}^{\rm U}$ is not provided in prior, i.e., the number of classification categories is unknown. Second, the ID indexes obtained from the classification of $\mathcal{T}$ and $\mathcal{S}$ are unaligned. To address the above problems, we propose a clustering-based identification consistency estimation strategy.
@@ III-D Implementation details
which are controlled by a learnable weight vector to determine the ratio among them. The student network is first trained with the supervised losses $\mathcal{L}_{\rm cse},\mathcal{L}_{\rm tri}$ and the domain discriminative loss $\mathcal{L}^{\rm S}$. For the training using the mean teacher strategy, the weights of the teacher network are initialized with those of the student network. The EMA method is used to update the weights of the teacher network in each training iteration. Then, the consistency loss $\mathcal{L}^{\mathrm{C}}_{1},\mathcal{L}^{\mathrm{C}}_{2}$ can be applied. Note that, under the standard UDA setting, in the training stage, both synthetic data with ID annotations and and real data (training set) without any annotation are used for training the network.
Inference pipeline. Following the classical mean-teacher schema, during inference, only the trained student network is used for testing and evaluation on the real data (testing set).
@@
We use the K-means algorithm to partition the video-level features into $M$ clusters, where $M$ is set as 16. We use 4 $\times$ NVIDIA RTX 2080ti for model training and apply the Adam [44] algorithm for optimization. During the testing phase, given a query, the trained student network retrieves a ranked list from the gallery.
IV Experimental Results
IV-A Setup
@@
Synthetic video Re-ID dataset. Most previous works like RandPerson [35], UnrealPerson [8], and ClonePerson [36] for synthetic Re-ID data generation only focus on the images. A recent work [41] uses the highly realistic video game Grand Theft Auto (GTA) V to construct a large-scale dataset for clothes-changing video person Re-ID. The videos were collected in a surveillance scenario, where 10 monitoring cameras were installed across five different scenes. We notice that this dataset mainly focuses on the clothes-changing challenge in the Re-ID problem, in which the persons have various (2-37) suits. Based on this dataset, we built a Synthetic-domain-adaptive Video person Re-ID dataset, namely SVReID. Specifically, we first only consider the non-clothing change scenario. From the training set in  [41], we extract a subset by selecting the highest number of sequences for each person without clothing variation and get SVReID including 1,164 sequences from 333 identities. We also consider the clothes-changing problem in some real-world video Re-ID datasets, e.g., CCVID [12]. This way, we further extend SVReID by including the clothes-changing videos, i.e., we select the top-8 highest number of sequences with clothing variations and get a total of 3,187 sequences, which form SVReID (w cc). Note that this work mainly focuses on the cross-domain problem, thus we keep the clothing scenario of the training data the same as the corresponding testing dataset. This prevents the model from excessively focusing on irrelevant data, and we can compare the performance of different methods under the same scenario. This synthetic dataset has two main advantages. Reality: The videos generated by GTA-V are very realistic, as shown in Fig. 3, the involved persons have various gaits, silhouettes even changed clothes. Diversity: It offers a wider range of various conditions, such as the camera view angles, illumination, and occlusion/clutter scenarios, which can not be easily achieved in real data.
Real-world video Re-ID datasets. In this work, we evaluate the effectiveness of our domain-adaptive Re-ID method on four prominent real-world video datasets, i.e., MARS [45], iLIDS-VID [46], PRID-2011 [47], LS-VID [48], and CCVID [12].
@@
$\bullet$ CCVID is a Re-ID dataset CCVID derived from a gait dataset FVG [49]. The characteristic of it is with clothing change setting. CCVID contains 2,856 video sequences from 226 identities, with an average of 2 outfits per person, and an average of 121 frames per sequence.
@@
TABLE I: Comparison with the state-of-the-art Re-ID methods on four pairs of cross-domain datasets, i.e., from the synthetic SVReID to the real datasets including Mars, iLIDS-VID, PRID, and LS-VID. The values with bold and underline mean the best and second best performance, respectively. (%)